22-42=4(x-1) solve for x
Answer by nyc_function(2741)

22-42=4(x-1)
22 - 42 = 4x - 4
-20 = 4x - 4
-20 + 4 = 4x
-16 = 4x
-16/4 = x
-4 = x
THE PROVE:
Let x = -4
22 - 42 = 4((-4)-1)
-20 = 4(-5)
-20 = -20
The answer is x = -4